OpenStudy (anonymous):

how would i simplify this and write it in standard form: X = 288/ (x+5)(x-4)?? please help

OpenStudy (anonymous):

please helP?

OpenStudy (anonymous):

would it be x=288 / (x+1)

OpenStudy (mathmate):

It is not clear the second set of parentheses belong to the numerator or denominator. I prefer to see parentheses around EACH of numerator and denominator.

OpenStudy (anonymous):

okay! sorry : (288) / (x+5)(x-4)

OpenStudy (mathmate):

It is also not clear if X is the same as x, or they represent different things.

OpenStudy (anonymous):

Different things.

OpenStudy (anonymous):

Imean same sorry, I am trying to do this math problem and there are multiple steps. I will rewrite the equation Thank you for your patience.

OpenStudy (mathmate):

As it is, it means \[\frac{288}{x+5}(x-4)\] and we usually write [(288) / (x+5)](x-4) to make sure there is no ambiguity. If you mean\[\frac{288}{(x+5)(x-4)}\] we would understand if you wrote (288) / [(x+5)(x-4)]

OpenStudy (anonymous):

X = 288/ (x-5)(x+4)

OpenStudy (anonymous):

That'swhat i Meant (288) / (x-5)(x+4)

OpenStudy (mathmate):

There isn't much to simplify either way, since X is a different entity from x! The way it was written, we multiply/divide from left to right, so X=(288) / (x-5)(x+4) = (288/(x-5)) * (x+4) =288(x-4) / (x-5)

OpenStudy (anonymous):

Well what if x meant the same thing?

OpenStudy (anonymous):

Would it be (288) / (x-1)

OpenStudy (anonymous):

sorry for all the confusion but thank you for helping me

OpenStudy (mathmate):

No it is not. Please check the question. Sorry for not being of much help!